OpenStudy (anonymous):

An equation is shown below: 2x + y = 3 Part A: Explain how you will show all of the solutions that satisfy this equation. Part B: Determine three different solutions for this equation. Part C: Write an equation that can be paired with the given equation in order to form a system of equations that is inconsistent.

OpenStudy (anonymous):

@undeadknight26

undeadknight26 (undeadknight26):

http://openstudy.com/study#/updates/52d89218e4b05ee054379135

OpenStudy (anonymous):

>.<

OpenStudy (anonymous):

i already looked at that, i don't understand it

undeadknight26 (undeadknight26):

:O ._.

undeadknight26 (undeadknight26):

Part A: To show "all" solutions, you could solve for y giving you y = -2x + 3 This essentially saying that, in 2D space (cartesian plane), any point with the coordinates (x, -2x+3) solves that solution. We can show this solution as with a graph. In this case, all the points to your solution form a line with a y-intercept of 3 (implies x= 0) and a slope of -2. B. Choose any three points on that line and you get three solutions! Or just plug in 3 different x values and solve for y. Ex: Let x = 4 y = -2(4) + 3 = -8+3 = -5 (4,-5) C. A system of equations is "inconsistent" if no solutions exist. Graphically, this means that two lines are not intersecting. Can you draw a line that does not intersect that graph of our above equatioN?

OpenStudy (anonymous):

So part A for that gives the answer, so i gotta do B and C now? @undeadknight26

OpenStudy (anonymous):

hello?

undeadknight26 (undeadknight26):

Nope thats all three